40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 368     Entry time: Mon Sep 20 20:57:40 2010
Author: tara 
Type: DailyProgress 
Category: PMC 
Subject: a part of PMC servo is verified 

 

I measured PMC's TF between MIXER OUT and PCMON, the measurement agrees well with LISO model.

 

The source is 10mV swept sine. It is sent through FP2. MIXER OUT and PCMON are connected to chA and chB respectively.

The measurement was done twice, with loop closed (PMC is locked to the laser )and loop opened (PMC's enable switch is off.)

The data from LISO model is added by 30dB to match the gain slider.

 

The result from closed loop looks weird at low frequency, I'm not sure why so I'm reading Application note 243 to find out about coherence.

 

I did this to make sure that the schematic and the board are matching, all parts in the schematic are labeled correctly.

Please ignore fig1, I did not have phase plot in it.

Attachment 1: TFcompare.png  21 kB  | Hide | Hide all
TFcompare.png
Attachment 2: TFcompare.png  21 kB  Uploaded Mon Sep 20 22:11:51 2010  | Hide | Hide all
TFcompare.png
ELOG V3.1.3-